A high schools parent teacher organization has raised $1800 to help pay for a trip to an amusement park. They ask that there be one adult for every five students attending. Adult tickets cost $45 and student tickets cost $30. If the group wants to take 50 students, how much will each student need to pay so the adults agreeing to chaperone pay nothing
Answer by macston(5194)

.
x=number of chaperones needed.
x/50 students=1 adult/5students
x/10=1 adult
x=10 adults
To pay for 50 students and 10 adults:
50($30)+10($45)=$1500+$450=$1950
They have collected $1800, so they are $150 short.
$150/50students=$3/student
ANSWER: Each student must pay $3.00.